# RELIEF: Reinforcement Learning Empowered Graph Feature Prompt Tuning

### v2

@@ Abstract.
Graph neural networks, prompt tuning, reinforcement learning, hybrid action space, few-shot learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
In recent years, Graph Neural Networks (GNNs) have been applied in diverse domains, including knowledge graphs (Liang et al., 2023), social networks (Cao et al., 2021), and recommender systems (Jiang et al., 2023), due to their powerful expressivity in capturing complex relationships in real-world data. However, traditional supervised learning on graphs struggles with limited labeled data and poor generalization to out-of-distribution samples (Yehudai et al., 2021). To overcome these challenges, significant efforts have focused on pre-trained GNN models (Wu et al., 2021; Xie et al., 2022), which are first trained on label-free graphs in a self-supervised manner to learn intrinsic properties, followed by fine-tuning on downstream tasks.
Despite advancements, the “pre-train, fine-tune” paradigm still exhibits drawbacks. The misalignment between pre-training and downstream objectives can lead to sub-optimal performance, known as negative transfer (Han et al., 2021). Moreover, under few-shot settings, pre-trained models are prone to overfitting, leading to catastrophic forgetting (Liu et al., 2021), which compromises their generalization ability. To tackle the issues, prompt learning, which has shown remarkable success in Natural Language Processing (NLP) (Brown et al., 2020; Liu et al., 2023b; Lester et al., 2021) and Computer Vision (CV) (Jia et al., 2022; Kirillov et al., 2023; Yao et al., 2023), emerges as a promising approach. By designing informative prompts to manipulate input data, prompt tuning aligns objectives between pretext and downstream tasks, thereby augmenting the potential of pre-trained models.
Naturally, researchers have extended the “pre-train, prompt” framework to graph domain (Sun et al., 2023b). Existing approaches are broadly divided into two categories based on whether they depend on pre-training strategies. For methods reliant on pre-training strategies, GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) unify pretext and downstream tasks as edge prediction. All in One (Sun et al., 2023a) introduces a prompt graph to match the graph-level contrastive learning used in pre-training. SGL-PT (Zhu et al., 2023) aligns the downstream task with masked node prediction using contrastive and generative learning objectives. However, these methods may fail if a GNN model is pre-trained with multiple self-supervised techniques through multi-task learning to enhance model’s capabilities, rather than solely simple link prediction.
In contrast, prompting methods agnostic to pre-training strategies offer greater compatibility. GPF (Fang et al., 2024), a pioneering method in this line, introduces graph feature-based prompt tuning, demonstrating that adding a learnable uniform feature prompt vector to each node is theoretically equivalent to any graph manipulation, without restricting pre-training strategies. GPF-plus (Fang et al., 2024), a variant of GPF, provides more flexible adaptation by inserting node-specific prompted features. SUPT (Lee et al., 2024), a subsequent study, innovated by assigning feature prompts at the subgraph level, captures intricate contextual nuances of each subgraph, leading to a decent performance improvement. Thus, feature-based prompting forms the basis of our paper due to its generality and effectiveness.
@@
We propose that given an adequately pre-trained GNN model, judiciously inserting minimal prompts as conditioning signals may suffice for downstream tasks. In the realm of feature prompting, “minimal” can be translated as an appropriate amount and magnitude of prompts. Conversely, incorporating prompts to an excessive number of nodes increases learnable parameters, raising the risk of overfitting (Li and Liang, 2021), particularly with few-shot downstream data. Moreover, since prompts are directly added to node features, large-magnitude prompts might overwhelm the original input’s feature space, causing pre-trained GNNs to perceive prompted graphs as unfamiliar, leading to a decline in transfer performance.
Hence, the key lies in strategically incorporating necessary and lightweight prompts into the original graph, allowing a GNN model to fully leverage its learned knowledge during pre-training without being overly disrupted, and to better generalize to downstream tasks through flexible feature prompts.
The choice of prompting on certain nodes and attaching specific feature prompts leads to a combinatorial optimization problem. To tackle this, we turn to Reinforcement Learning (RL), which effectively searches for heuristic strategies by training an agent (Mazyavkina et al., 2021), and propose a REinforcement LearnIng Empowered graph Feature prompting method, named RELIEF. We formulate the process of inserting prompts as a sequential decision-making problem. At each step, the RL agent: 1) selects which node to prompt (discrete action), and 2) determines the prompt content, i.e., specific values of the prompt vector (continuous action). Therefore, an RL algorithm capable of handling the discrete-continuous hybrid action space (Fan et al., 2019; Masson et al., 2016; Xiong et al., 2018) is employed. The prompted graph is then evaluated by the pre-trained GNN model. At the subsequent step, a new prompt generated by the RL agent is added to the previous prompted graph. Prompt addition and performance evaluation are repeated until a maximum step is reached. Our goal is to maximize the expected cumulative performance gain on the downstream task. Techniques like policy generalization (Ghosh et al., 2021; Wang et al., 2020; Raileanu and Fergus, 2021) are integrated to ensure training stability and effectiveness. Furthermore, we have designed two metrics: prompt coverage ratio and average prompt magnitude to quantify the extent to which the prompt affects the original input.
Our contributions are summarized as follows:
@@ 2. Preliminaries
Most RL algorithms operate within either discrete or continuous action spaces  (Sutton and Barto, 2018). In our scenario, however, we deal with a hybrid action space  (Masson et al., 2016; Xiong et al., 2018), involving selecting nodes (discrete) and specifying feature prompts (continuous). Formally, discrete actions are chosen from a finite set $\mathcal{A}_{d}=\{a_{1},a_{2},\dots,a_{n}\}$, each corresponding to a real-valued continuous action $z\in\mathbb{R}^{D}$, where $D$ is the dimension of the continuous action space $\mathcal{A}_{c}=\{z_{1},z_{2},\dots,z_{n}\}$. H-PPO  (Fan et al., 2019) is an RL algorithm tailored for hybrid action spaces based on actor-critic architecture. It employs two parallel actor networks for discrete and continuous actions, respectively, alongside a single critic network for value estimation. H-PPO extends Proximal Policy Optimization (PPO)  (Schulman et al., 2017) by separately optimizing the two actors with an entropy-regularized PPO surrogate objective $\mathcal{L}^{\text{PPO}}$ to maximize expected cumulative discounted reward, while the critic minimizes Mean Squared Error (MSE) loss $\mathcal{L}^{\text{Critic}}$ to improve value estimation accuracy. For a detailed introduction to PPO and H-PPO, please refer to Appendix  E.
@@ 3.4. Policy Generalization
To sum up, the policy network, enhanced by policy generalization techniques, now comprises $l$ discrete actors, $l$ continuous actors and a single critic. During training, discrete and continuous actors within each pair are independently updated using Eq. (4) and (6), respectively, followed by the critic update using $\mathcal{L}^{\text{Critic}}$. Each pair of actors is sequentially updated. The joint policies $\pi_{d,\,J}$ and $\pi_{c,\,J}$ are applied to incorporate prompt features when training the projection head or testing at the evaluation stage. The pseudo-code for RELIEF can be found in Appendix F. Additionally, a detailed complexity analysis is in Appendix G.
@@
3.5. Metrics for Quantifying Prompts Impact
Prompt Coverage Ratio.
Average Prompt Magnitude.
@@ 5.1. Few-shot Graph Classification
We use a 5-layer GIN (Xu et al., 2018) as the base architecture for the GNN model, which is pre-trained on chemistry datasets (Hu et al., 2019) and prompt tuned on molecule properties prediction benchmark from MoleculeNet  (Wu et al., 2018) as downstream tasks, following the setup in existing feature prompting works (Fang et al., 2024; Lee et al., 2024). Detailed descriptions of these datasets are provided in Appendix B.1.
@@
To demonstrate the generality of our RELIEF, we pre-train the GIN model with four common strategies at both node and graph levels: Deep Graph Infomax (Infomax) (Veličković et al., 2018), Attribute Masking (AttrMasking) (Hu et al., 2019), Context Prediction (ContextPred) (Hu et al., 2019), and Graph Contrastive Learning (GCL) (You et al., 2020). Detailed descriptions of these methods are available in Appendix B.2.
@@
We compare RELIEF with the following tuning approaches, all of which involve adjusting the projection head: additionally, Fine-Tuning (FT) tunes the parameters of the pre-trained GNN model; feature prompt tuning methods, including GPF, GPF-plus (Fang et al., 2024), SUPTsoft and SUPThard (Lee et al., 2024), tune the parameters of learnable feature prompts with the GNN model fixed; other prompt-based method, All in One (Sun et al., 2023a), uses a frozen GNN model specifically pre-trained by GCL, adapting learnable graph prompts, is included as it strictly aligns with our experimental setting.
@@
Since molecular predictions are multi-label binary classifications, we use binary cross-entropy loss for the reward function and evaluate by averaging the ROC-AUC of each label. We construct few-shot scenarios with 50 samples for prompt tuning. We leverage early stopping based on validation set performance, unlike previous works that report results from the last epoch. Five rounds of experiments are conducted with different random seeds. The number of sub-policies $l$ is set to 3, with penalty factors $\alpha_{d}$ selected from {1e0, 1e1, 1e3, 1e5} and $\alpha_{c}$ set to 0.1. The maximum prompt scale per step $z_{\text{max}}$ is selected from $\{0.1,0.5,1.0\}$. The projection head is chosen from $\{1,2,3\}$-layer MLPs and trained $q$ times per epoch, ranging from $\{1,2,3\}$. Detailed data splitting and hyper-parameter settings are provided in Appendix B.3. All the experiments were conducted on an NVIDIA A800 PCIe 80GB GPU.
@@
Based on the results in Table 1, our RELIEF achieves superior graph classification performance in few-shot settings, surpassing baselines in 28/32 tasks, with an average improvement of 1.64% over fine-tuning and 1.04% over the runner-up. Notably, RELIEF is the only method consistently outperforming fine-tuning across all tasks. Moreover, we observe that in All in One, inserting prompt graphs with fewer nodes yields better results. This suggests that compact prompt structures may enhance prompt tuning, indirectly supporting our motivation.
We illustrate the tuning process of RELIEF on the BACE dataset pre-trained with Infomax in Figure 3. The ROC-AUC curve demonstrates a smooth and monotonically increasing trend, indicating the stability of the training process. The reward curve and distributions show reward improvement as the epochs progress, confirming the effectiveness of RELIEF’s agent in enhancing downstream task performance. Detailed tuning process analysis is in Appendix B.4.
Furthermore, we measure the impact of prompts on input graphs in terms of PCR and APM, and further multiply them to express the overall impact (OV.). Results are averaged across all tasks, as outlined in Table 2. RELIEF exhibits the smallest PCR, APM and overall impact compared to baselines. Interestingly, SUPThard also prompts on specific nodes with a PCR of 0.65, as each basis prompt vector affects only the top-ranked nodes, where the top-rank ratio is a hyper-parameter, making it a rule-based node selection method. In contrast, RELIEF does not rely on priori information, such as the usage of basis vectors or a top-rank ratio, thereby offers a highly flexible prompting policy based on the actors. Notably, the flexibility of graph manipulations is crucial for prompting performance (Sun et al., 2023a, b), contributing to RELIEF’s superior results.
@@ 5.2. Data Efficiency
To evaluate the data efficiency of feature prompting methods, we incrementally add 5% of the training samples until the average ROC-AUC over five random seeds surpasses full-shot fine-tuning performance. We tune hyper-parameters after each data increment to optimize performance. We report the minimum proportion of data required for each method to exceed fine-tuning on five datasets pre-trained with Infomax and ContextPred strategies.
Table 3 shows that RELIEF requires the least data for all tasks, while other baselines, more or less, fail to surpass full-shot fine-tuning even with the entire dataset. Figure 4 illustrates the performance trends on BBBP dataset pre-trained with Infomax, which demonstrates that RELIEF’s overall performance presents the most noticeable improvement as data scales up, whereas other methods fluctuate and ultimately fail to exceed fine-tuning.
We attribute RELIEF’s superior data efficiency to its RL-based learning paradigm, which posses two advantages. Firstly, it reduces learning difficulty by optimizing one feature prompt at a time, allowing the agent to focus on a single state-to-action mapping and avoid complex node interactions that baseline methods face when optimizing entire feature prompts at once. Secondly, the step-by-step insertion and evaluation of prompts expose the agent to various state patterns and loss values, acting as data augmentation and providing additional supervision signals in few-shot settings.
5.3. Additional Experiments
Due to space constraints, we relocate additional experiments to the Appendices. Here, we provide an overview to highlight their critical conclusions.
Node Classification (Appendix  A). We extend RELIEF to node-level tasks by adding prompts to the $k$-hop subgraph induced by the target node (Liu et al., 2023a; Sun et al., 2023a). Experiments demonstrate RELIEF’s top-tier performance across multiple datasets compared to baselines. A case study is conducted to reveal that RELIEF can achieve its full potential when given adequately pre-trained GNN models.
Ablation Study (Appendix  C). We compare RELIEF against its three variants: random discrete policy, random continuous policy, and linear probing (solely training the projection head) (Kumar et al., 2022) on node and graph-level tasks. Results indicate significant performance drops when either policy is deactivated or feature prompting components are discarded.
Parameter Analysis (Appendix  D). We examine the sensitivity of key hyper-parameters of RELIEF including the maximum feature prompt scale per step $z_{\text{max}}$, number of sub-policies $l$, and penalty factors $\alpha_{d}$ and $\alpha_{c}$. We also validate the necessity of using policy generalization techniques. The findings confirm RELIEF’s robust performance across varied hyper-parameter settings.
5.4. Why RELIEF works?
Lightweight. As the case study analyzed in Appendix A.6, RELIEF tends to preserve pre-trained graph knowledge rather than overwhelm it. Larger-scale prompts cause the GNN model to generate unfamiliar representations, disrupting coordination between representations and the projection head, potentially leading to negative rewards. These punishments are applied to the actor, urging it to provide prompts that avoid excessive disturbance to the initial input through policy updates.
6. Conclusion
In this paper, inspired by the marginal effect of increasing prompt token length on performance improvement in LLMs, we propose RELIEF, an RL empowered graph feature prompt tuning method. RELIEF enhances the downstream performance of pre-trained GNN models by incorporating only necessary and lightweight feature prompts to original graphs. To tackle the combinatorial optimization problem, we employ an RL algorithm designed for hybrid action spaces and integrated policy generalization techniques. Extensive experiments on graph and node-level tasks demonstrate RELIEF’s effectiveness and data efficiency compared to fine-tuning and other prompt-based methods.
Acknowledgements.
Appendix A Few-Shot Node Classification
A.1. Extend to Node-Level Tasks
Given that feature prompting methods are originally designed to address graph-level tasks (Fang et al., 2024), adapting them to node-level tasks requires careful consideration.

### v3

@@ Abstract.
Graph Neural Networks; Prompt Tuning; Reinforcement Learning; Hybrid Action Space; Few-shot Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709252isbn: 979-8-4007-1245-6/25/08ccs: Information systems Data miningccs: Computing methodologies Reinforcement learningccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
In recent years, Graph Neural Networks (GNNs) have been applied in diverse domains, including knowledge graphs (Liang et al., 2023), social networks (Cao et al., 2021), and recommender systems (Jiang et al., 2023), due to their powerful expressivity in capturing complex relationships in real-world data. To facilitate robust graph learning under few-shot and out-of-distribution settings, significant efforts have focused on pre-training GNN models (Xie et al., 2022). Despite advancements, the “pre-train, fine-tune” paradigm still exhibits drawbacks. The misalignment between pre-training and downstream objectives can lead to sub-optimal performance (Han et al., 2021). Moreover, pre-trained models are prone to overfitting given rather limited samples, leading to catastrophic forgetting (Liu et al., 2021), which compromises their generalization ability.
To tackle the issues, prompt learning, which has shown remarkable success in Natural Language Processing (NLP) (Liu et al., 2023b; Lester et al., 2021) and Computer Vision (CV) (Jia et al., 2022; Yao et al., 2023), emerges as a promising approach. By designing informative prompts to manipulate input data, prompt tuning aligns objectives between pretext and downstream tasks, thereby augmenting the potential of pre-trained models.
Naturally, researchers have extended the “pre-train, prompt” framework to graph domain (Sun et al., 2023b; Gao et al., 2024; Wang et al., 2024; Jin et al., 2024; Zi et al., 2024). Existing methods are broadly divided into two categories based on dependence on pre-training strategies. For methods reliant on pre-training strategies, GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) unify pretext and downstream tasks as edge prediction. All in One (Sun et al., 2023a) introduces a prompt graph to match the graph-level contrastive learning used in pre-training. However, these methods may fail if the pre-training and prompt tuning objectives of the GNN model are misaligned. In contrast, prompting methods agnostic to pre-training strategies offer greater compatibility. GPF (Fang et al., 2024) introduces graph feature-based prompt tuning, demonstrating that adding a learnable uniform feature prompt vector to each node is theoretically equivalent to any graph manipulation, without restricting pre-training strategies. GPF-plus (Fang et al., 2024) and SUPT (Lee et al., 2024) further provide more flexible adaptation by inserting node or subgraph-specific prompted features, respectively, capturing intricate topological nuances of graph, leading to a decent performance improvement. Thus, feature-based prompting forms the basis of our paper due to its generality and effectiveness.
@@
We propose that given a pre-trained GNN model, judiciously inserting minimal prompts as conditioning signals may suffice for downstream tasks. In the realm of feature prompting, “minimal” can be translated as an appropriate amount and magnitude of prompts. Conversely, since prompts are directly added to node features, incorporating large-magnitude prompts to an excessive number of nodes might overwhelm the original input’s feature space, causing pre-trained GNNs to perceive prompted graphs as unfamiliar, leading to a decline in transfer performance. Hence, the key lies in strategically incorporating necessary and lightweight prompts into the original graph, allowing a GNN model to fully leverage its learned knowledge during pre-training without being overly disrupted, and to better generalize to downstream tasks through flexible feature prompts.
The choice of prompting on certain nodes and attaching specific prompt contents leads to a combinatorial optimization problem. To tackle this, we turn to Reinforcement Learning (RL), which effectively searches for heuristic strategies by training an agent (Mazyavkina et al., 2021), and propose a REinforcement LearnIng Empowered graph Feature prompting method, named RELIEF. We formulate the process of inserting prompts as a sequential decision-making problem. At each step, the RL agent: 1) selects which node to prompt (discrete action), and 2) determines the prompt content, i.e., specific values of the prompt vector (continuous action). Therefore, an RL algorithm capable of handling the discrete-continuous hybrid action space (Fan et al., 2019; Masson et al., 2016) is employed. The prompted graph is then evaluated by the pre-trained GNN model. At the subsequent step, a new prompt generated by the RL agent is added to the previous prompted graph. Prompt addition and performance evaluation are iteratively performed to maximize the expected cumulative performance gain on the downstream task. Techniques like policy generalization (Ghosh et al., 2021; Wang et al., 2020; Raileanu and Fergus, 2021) are integrated to ensure training stability and effectiveness. Our contributions are summarized as follows:
@@ 2. Preliminaries
Most RL algorithms operate within either discrete or continuous action spaces  (Sutton and Barto, 2018). In our scenario, however, we deal with a hybrid action space  (Masson et al., 2016), involving selecting nodes (discrete) and specifying feature prompts (continuous). Formally, discrete actions are chosen from a finite set $\mathcal{A}_{d}=\{a_{1},a_{2},\dots,a_{n}\}$, each corresponding to a real-valued continuous action $z\in\mathbb{R}^{D}$, where $D$ is the dimension of the continuous action space $\mathcal{A}_{c}=\{z_{1},z_{2},\dots,z_{n}\}$. H-PPO  (Fan et al., 2019) is an RL algorithm tailored for hybrid action spaces based on actor-critic architecture. It employs two parallel actor networks for discrete and continuous actions, respectively, alongside a single critic network for value estimation. H-PPO extends Proximal Policy Optimization (PPO)  (Schulman et al., 2017) by separately optimizing the two actors with an entropy-regularized PPO surrogate objective $\mathcal{L}^{\text{PPO}}$ to maximize expected cumulative discounted reward, while the critic minimizes Mean Squared Error (MSE) loss $\mathcal{L}^{\text{Critic}}$ to improve value estimation accuracy.
@@ 3.4. Policy Generalization
To sum up, the policy network, enhanced by policy generalization techniques, now comprises $l$ discrete actors, $l$ continuous actors and a single critic. During training, discrete and continuous actors within each pair are independently updated using Eq. (4) and (6), respectively, followed by the critic update using $\mathcal{L}^{\text{Critic}}$. Each pair of actors is sequentially updated. The joint policies $\pi_{d,\,J}$ and $\pi_{c,\,J}$ are applied to incorporate prompt features when training the projection head or testing at the evaluation stage.
3.5. Metrics for Quantifying Prompts Impact
Prompt Coverage Ratio.
Average Prompt Magnitude.
@@ 5.1. Few-shot Graph Classification
We use a 5-layer GIN (Xu et al., 2018) as the base architecture for the GNN model, which is pre-trained on chemistry datasets (Hu et al., 2019) and prompt tuned on molecule properties prediction benchmark from MoleculeNet  (Wu et al., 2018) as downstream tasks. Descriptions of these datasets are provided in Appendix A.1.
@@
To demonstrate the generality of our RELIEF, we pre-train the GIN model with four common strategies at both node and graph levels: Deep Graph Infomax (Infomax) (Veličković et al., 2018), Attribute Masking (AttrMasking) (Hu et al., 2019), Context Prediction (ContextPred) (Hu et al., 2019), and Graph Contrastive Learning (GCL) (You et al., 2020). Descriptions of these methods are available in Appendix A.2.
@@
We compare RELIEF with the following tuning approaches, all of which involve adjusting the projection head. Besides, Fine-Tuning (FT) tunes the parameters of the pre-trained GNN model; feature prompt tuning methods, including GPF, GPF-plus (Fang et al., 2024), SUPTsoft and SUPThard (Lee et al., 2024), tune the parameters of learnable feature prompts with the GNN model fixed; other prompt-based method, All in One (Sun et al., 2023a), uses a frozen GNN model specifically pre-trained by GCL, adapting learnable graph prompts, is included as it strictly aligns with our experimental setting.
@@
Since molecular predictions are multi-label binary classifications, we use binary cross-entropy loss as the reward function and evaluate by averaging the ROC-AUC of each label. We construct few-shot scenarios with 50 samples for prompt tuning over five random seeds. The number of sub-policies $l$ is set to 3, with penalty factors $\alpha_{d}$ selected from {1e0, 1e1, 1e3, 1e5} and $\alpha_{c}$ set to 0.1. The maximum prompt scale per step $z_{\text{max}}$ is selected from $\{0.1,0.5,1.0\}$. The projection head is chosen from $\{1,2,3\}$-layer MLPs and trained $q$ times per epoch, ranging from $\{1,2,3\}$. Detailed data splitting and hyper-parameter settings are provided in Appendix A.3. All the experiments were conducted on an NVIDIA A800 PCIe 80GB GPU.
@@
Based on the results in Table 1, RELIEF achieves superior graph classification performance in few-shot settings, surpassing baselines in 28/32 tasks, with an average improvement of 1.64% over fine-tuning and 1.04% over the runner-up. Notably, RELIEF is the only method consistently surpassing fine-tuning across all tasks. Moreover, we observe that All in One yields better results by inserting prompt graphs with fewer nodes, suggesting that compact prompt structures may enhance prompt tuning.
We illustrate the tuning process of RELIEF on the BACE dataset pre-trained with Infomax in Figure 3. The ROC-AUC curve demonstrates a monotonically increasing trend, indicating the stability of the training process. The reward curve and distributions show reward improvement as the epochs progress, confirming the effectiveness of RELIEF in enhancing downstream task performance.
Furthermore, we measure the impact of prompts on input graphs in terms of PCR and APM, and further multiply them to express the overall impact (OV.). Results are averaged across all tasks, as outlined in Table 2. RELIEF exhibits the smallest PCR, APM and overall impact compared to baselines. Interestingly, SUPThard also prompts on specific nodes with a PCR of 0.65, as each basis prompt vector affects only the top-ranked nodes, where the top-rank ratio is a hyper-parameter. In contrast, RELIEF does not rely on priori information, such as the usage of basis vectors or a top-rank ratio, offering a highly flexible prompting policy based on the actors. Notably, the flexibility of graph manipulations is crucial for prompting performance (Sun et al., 2023a, b), contributing to RELIEF’s superior results.
[…]
Figure 3 Fully described in the text.
@@ 5.2. Data Efficiency
To evaluate the data efficiency of feature prompting methods, we incrementally add 5% of the training samples until the average ROC-AUC over five random seeds surpasses full-shot fine-tuning. We report the minimum proportion of data required for each method on five datasets pre-trained with Infomax and ContextPred strategies.
Table 3 shows that RELIEF requires the least data for all tasks, while other baselines, more or less, fail to surpass full-shot fine-tuning even with the entire dataset. Figure 4 illustrates the performance trends on BBBP pre-trained with Infomax, which demonstrates that RELIEF achieves the most noticeable improvement as data scales up, whereas other methods fail to exceed fine-tuning.
We attribute RELIEF’s superior data efficiency to its RL-based learning paradigm. Firstly, it reduces learning difficulty by optimizing one feature prompt at a time, allowing the agent to focus on a single state-to-action mapping instead of optimizing entire feature prompts at once. Secondly, the step-by-step insertion and evaluation of prompts expose the agent to various state patterns and loss values, acting as data augmentation in few-shot settings.
@@
5.3. Few-shot Node Classification
Extend to Node-Level Tasks.